OpenStudy (anonymous):

Nakim simplified 3(sqrt)2x + x(sqrt)8x - 5(sqrt)18x and got -10(sqrt)2x explain where she went wrong and show your work?

OpenStudy (campbell_st):

\[3\sqrt{2x} + x \sqrt{8x} - 5\sqrt{18x} -10\sqrt{2x} = 3\sqrt{2x} - x \times 2\sqrt{2x} - 10 \times 3\sqrt{2x}\] each term has been expressed in terms of \[\sqrt{2x}\] so collect like terms \[(x -12)\sqrt{2x}\]

OpenStudy (campbell_st):

oops \[(2x - 12)\sqrt{2x}\]

jhonyy9 (jhonyy9):

3sqrt2x +xsqrt8x -5sqrt18x =3sqrt2x +2xsqrt2x -15sqrt2x =2xsqrt2x -12sqrt2x = =(2x-12)sqrt2x - so is wrong when was subtracted 2xsqrt2x from 12sqrt2x what not is possible

OpenStudy (anonymous):

thank you

jhonyy9 (jhonyy9):

yw good luck bye